# The kernel of formal polylogarithms

### v1

@@ 1 Introduction
One variable polylogarithmic functions (polylogs) were considered by Bernoulli and Leibniz at the end of the 17th century, and then by Spence in the 19th century. Polylogarithms in $n$ variables were defined by Goncharov [7]. They are given by power series
@@
Acknowledgments. This paper is part of the proceedings of the working group held by F. Brown and the authors at the University of Geneva from November 2023 to June 2024. We are grateful to B. Enriquez and H. Furusho for very interesting discussions.
@@ 3.2 Modules and actions
From now on, as will be the case in subsequent parts of this note, we assume the Lie algebra $\mathfrak{g}$ and the $\mathfrak{g}$-module $M$ to be graded with finite dimensional graded components. For a module with a cyclic vector $(M,v)$, the (graded) dual module $M^{*}$ can be realized as a subspace of $(U\mathfrak{g})^{*}$:
@@ 6.2 Computation of $\mathfrak{k}_{5}$ and $\mathfrak{pl}_{5}$
Recall that ${\rm ker}(\pi_{\beta})=\mathfrak{f}_{3}(X_{2\beta},X_{3\beta},X_{\beta\omega})$. Then,
${\rm ker}(\pi_{2})\cap{\rm ker}(\pi_{\beta})=[X_{2\beta}]\quad\text{and}\quad{\rm ker}(\pi_{3})\cap{\rm ker}(\pi_{\beta})=[X_{3\beta}],$
where $X_{2\beta}]$ and $X_{3\beta}]$ are Lie ideals in $\mathfrak{f}_{3}(X_{2\beta},X_{3\beta},X_{\beta\omega})$ generated by $X_{2\beta}$ and $X_{3\beta}$, respectively. Finally,
${\rm ker}(\pi_{2})\cap{\rm ker}(\pi_{3})\cap{\rm ker}(\pi_{\beta})=[X_{2\beta}]\cap[X_{3\beta}]=[[X_{2\beta}],[X_{3\beta}]]=\lambda,$
@@
Let $\psi\in\mathfrak{f}_{2}$ be an element of degree at least $3$ verifying the skew-symmetric relation $\psi(x,y)=-\psi(y,x)$. Then, Theorems $3.8$ and $3.15$ in [8] imply that the following two statements are equivalent:
  1. (i)
    $\psi\in\mathfrak{dmr}_{0},$ where $\mathfrak{dmr}_{0}$ is the double shuffle Lie algebra of Racinet [9];
  2. (ii)
    ${\rm pent}(\psi)\in[\lambda,\lambda]$, where ${\rm pent}(\psi)$ is the infinitesimal pentagon equation for $\psi$ (for more details, see [4]).
Observe that the condition (ii) is equivalent to the condition
${\rm pent}(\psi)\in{\rm ker}(\pi),$ (17)
relating the defect in the pentagon relation to the kernel of formal polylogs.

### v2

@@ 1 Introduction
One variable polylogarithmic functions (polylogs) were considered by Bernoulli and Leibniz at the end of the 17th century, and then by Spence in the 19th century. Polylogarithms in $n$ variables were defined by Goncharov [7]. They are given by the power series
@@
Acknowledgments. This paper is part of the proceedings of the working group held by F. Brown and the authors at the University of Geneva from November 2023 to June 2024. We are grateful to B. Enriquez, H. Furusho, and N. Markarian for very interesting discussions and for comments on the draft of this paper.
@@ 3.2 Modules and actions
From now on, as will be the case in subsequent parts of this note, we assume the Lie algebra $\mathfrak{g}$ and the $\mathfrak{g}$-module $M$ to be graded with finite-dimensional graded components. For a module with a cyclic vector $(M,v)$, the (graded) dual module $M^{*}$ can be realized as a subspace of $(U\mathfrak{g})^{*}$:
@@ 6.2 Computation of $\mathfrak{k}_{5}$ and $\mathfrak{pl}_{5}$
Recall that ${\rm ker}(\pi_{\beta})=\mathfrak{f}_{3}(X_{1\beta},X_{2\beta},X_{\beta\omega})$. Then,
${\rm ker}(\pi_{1})\cap{\rm ker}(\pi_{\beta})=[X_{1\beta}]\quad\text{and}\quad{\rm ker}(\pi_{2})\cap{\rm ker}(\pi_{\beta})=[X_{2\beta}],$
where $X_{1\beta}]$ and $X_{2\beta}]$ are Lie ideals in $\mathfrak{f}_{3}(X_{1\beta},X_{2\beta},X_{\beta\omega})$ generated by $X_{1\beta}$ and $X_{2\beta}$, respectively. Finally,
${\rm ker}(\pi_{1})\cap{\rm ker}(\pi_{2})\cap{\rm ker}(\pi_{\beta})=[X_{1\beta}]\cap[X_{2\beta}]=[[X_{1\beta}],[X_{2\beta}]]=\lambda,$
@@
Let $\psi\in\mathfrak{f}_{2}$ be an element of degree at least 3. Denote by ${\rm pent}(\psi)\in\mathfrak{p}_{5}$ the left hand side of the infinitesimal pentagon equation for $\psi$ (for more details, see [4]):
$\begin{array}[]{lll}{\rm pent}(\psi)&=&\psi(X_{\alpha 1},X_{12})++\psi(X_{\alpha 1}+X_{\alpha 2},X_{1\beta}+X_{2\beta})+\psi(X_{12},X_{2\beta})\\ &-&\psi(X_{\alpha 2}+X_{12},X_{2\beta})-\psi(X_{\alpha 1},X_{1\beta}+X_{12}).\end{array}$
The expression ${\rm pent}(\psi)$ belongs to the joint kernel of projections $\pi_{1},\pi_{2}$ and $\pi_{\beta}$. Hence, by Proposition 6.8, it belongs to the Lie ideal $\lambda$. B. Enriquez and H. Furusho suggested that the condition
${\rm pent}(\psi)\in[\lambda,\lambda]$ (17)
is equivalent to $\psi\in\mathfrak{dmr}_{0}$. Note that by Theorem 6.2 equation (17) is equivalent to the condition
${\rm pent}(\psi)\in{\rm ker}(\pi)$ (18)
stating that ${\rm pent}(\psi)$ belongs to the joint kernel of formal polylogs.
Remark 6.10.
Let $\psi\in\mathfrak{f}_{2}$ be an element of degree at least 33 verifying the skew-symmetric relation $\psi(x,y)=-\psi(y,x)$. Then, Theorems $3.8$ and $3.15$ in [8] imply the equivalence of the condition (18) to $\psi\in\mathfrak{dmr}_{0}$.
Remark 6.11.
N. Markarian has kindly informed us that he proved the following statement. Let $p_{i}\colon\pi_{1}^{un}(\mathcal{M}_{0,5})\to\pi_{1}^{un}(\mathcal{M}_{0,4})$ be the map forgetting one points, $\Lambda=\cap_{i=1,2,3}\ker p_{i}$ the joint kernel of three projections, and $f~\in~\pi_{1}^{un}(\mathcal{M}_{0,4})$ be symmetric element (that is, $f(x,y)=f^{-1}(y,x)$). In this case, ${\rm Pent}(f)\in\Lambda$, and coefficients of $f$ obey regularized double shuffle relations if and only if ${\rm Pent}(f)^{ab}\in\Lambda^{ab}$ vanishes. This statement is mentioned in the slides of Markarian’s talk [9] (also with the reference to Enriquez-Furusho), and it will be explained in his forthcoming paper “Multiplicative convolution and double shuffle relations”. The statement of Remark 6.10 resembles of its infinitesimal version.